# Automation of Quantum Dot Measurement Analysis via Explainable Machine Learning

### v4

@@ Abstract
The rapid development of quantum dot (QD) devices for quantum computing has necessitated more efficient and automated methods for device characterization and tuning. Many of the measurements acquired during the tuning process come in the form of images that need to be properly analyzed to guide the subsequent tuning steps. By design, features present in such images capture certain behaviors or states of the measured QD devices. When considered carefully, such features can aid the control and calibration of QD devices. An important example of such images are so-called triangle plots, which visually represent current flow and reveal characteristics important for QD device calibration. While image-based classification tools, such as convolutional neural networks (CNNs), can be used to verify whether a given measurement is good and thus warrants the initiation of the next phase of tuning, they do not provide any insights into how the device should be adjusted in the case of bad images. This is because CNNs sacrifice prediction and model intelligibility for high accuracy. To ameliorate this trade-off, a recent study introduced an image vectorization approach that relies on the Gabor wavelet transform [1]. Here we propose an alternative vectorization method that involves mathematical modeling of synthetic triangles to mimic the experimental data. Using explainable boosting machines, we show that this new method offers superior explainability of model prediction without sacrificing accuracy. This work demonstrates the feasibility and advantages of applying explainable machine learning techniques to the analysis of quantum dot measurements, paving the way for further advances in automated and transparent QD device tuning.
@@ 1 Introduction
In many machine learning (ML) applications, there has been a longstanding trade-off between the accuracy and interoperability of candidate models [2, 3, 4]. This is evident in the extreme example of deep neural networks (DNNs), which can offer excellent accuracy for many problems, often surpassing existing methods [5, 6, 7]. Yet the best-performing ML models are limited in their interpretability due to the number of inaccessible layers. Alternatively, simple techniques, such as linear models or decision trees, allow the user to fully comprehend the internal weights. However, these techniques often cannot model the complex relationships seen in modern datasets. For image data, there has been considerable progress toward finding a middle ground, typically through explaining complex models with surrogates such as the Local Interpretable Model-agnostic Explanations (LIME) [8] and Shapley [9].
@@
In contrast with many black-box ML models, explainable boosting machines (EBMs) are a glass-box method that enables the model to be directly interpretable as opposed to relying on surrogate explanations [10]. Specifically, EBMs extend generalized additive models to include pairwise interactions, which allows one to observe the relationship between features. EBMs often provide accuracy on par with many black-box models with the additional advantage of enhanced intelligibility, which makes them an appealing replacement for other models, especially in applications of consequence, such as medicine [11, 12, 13] or finance [14]. However, to date, EBMs have not been adapted to any other data type than tabular data.
@@
For spatial data, such as images, interpretability is more challenging. This is in part due to their composition by structures, which are highly correlated at multiple scales in the two-dimensional (2D) space. This feature is one of the main reasons why convolutional neural networks (CNNs) [15] and, more recently, vision transformers (ViTs) [16] have quickly become the dominant ML approach for many computer vision tasks. However, the black-box nature of CNNs and ViTs makes their use prohibitive in applications where a solid understanding of model predictions is necessary. While making CNNs and ViTs more interpretable has been a very active area of research, techniques proposed to date vary in utility in many applications, especially as the depth of the neural network increases [17, 18, 19].
@@
Our previous work developed a methodology that addresses some of these interpretability concerns for measurement data by combining image vectorization with EBMs. The possibility of using EBMs as models for image data poses numerous challenges, the principal of which is the mapping from images to a vector representation that could then be used directly with EBMs. In that work, we used the Gabor Wavelet transform in conjunction with a constrained optimization procedure to extract key image features from the data [20]. We also applied custom feature engineering to tailor this process to the particular dataset [1]. In both cases, we relied on domain knowledge and understanding of the physical systems under investigation to inform the feature extraction process. This was essential in ensuring that the resulting model produces human-agreeable interpretations. Here, we demonstrate that the same approach can be successfully applied to assist in the tune-up of accumulation mode Si/SixGe1-x quantum dot (QD) devices. We also propose an alternative image vectorization method involving the generation of synthetic data to approximate the experimentally acquired scans [21]. We then show that both methods result in a comparable performance, but the latter produces features that are easier to interpret.
@@
The paper is organized as follows: In section 2, we provide a brief overview of the scientific context of the problem. Section 2.1 describes the problem of tuning QD devices and introduces the concept of triangle plots. Data used for benchmarking is discussed in section 2.2. The two vectorization methods, the Gabor filterbank approach and the synthetic triangle fitting approach, are described in section 2.3 and section 2.5, respectively. Finally, the EBMs performance on experimental data vectorized using both methods, as well as using a hybrid approach combing Gabor filters with synthetic triangles, is presented in section 3. We conclude with a discussion of the future direction in section 4.
@@ 2 Background and methods
In this section, we introduce the scientific context of our problem and the data utilized in our experiments. We also present an overview of the signal processing methods used in the Gabor filterbank approach and the synthetic triangle fitting approach.
@@ 2.1 Quantum dot tuning problem and triangle plots
Arrays of QDs – interconnected islands of electrons confined in a semiconductor heterostructure with unique properties that allow them to act as artificial atoms – are a leading candidate for use as qubits, the fundamental information carriers in quantum computers [22]. We use a type of gate-defined QD device, as shown in figure 1(a), in which three layers of overlapping gates fabricated on top of a Si/SixGe1-x heterostructure are used to form and control QDs [23, 24]. Scaling these systems to large arrays suitable for quantum computations becomes hard, as with the number of QDs, the number of gates needed to control them grows, making the manual tuning process unfeasible. An autotuning framework incorporating ML tools was originally proposed and validated off-line in Ref. [25] using premeasured experimental scans capturing a large range of gate voltages and then deployed online (i.e., in situ) to tune a double QD in real-time in [26]. A detailed description of the tuning process is available in Ref. [27].
@@
Figure 1: (a) False-colored SEM micrograph of a typical Si/SixGe1-x heterostructure device. The red highlighted gates are the screening gates that are swept for the triangle plot. This device has three distinct current channels between the screening gates marked with white arrows. (b) An example of a good triangle plot with high visibility in the triangle region above the background and (c) a corresponding synthetic triangle plot. (c) An example of a bad triangle plot with little to no current in the triangle region and (d) a corresponding synthetic triangle plot.
@@
One sub-problem within the tuning procedure is to correctly determine the voltage placement of gates to allow the formation of isolated current channels inside the 2D electron gas formed at the intersection of the Si and SixGe1-x layers in the heterostructure, see figure 1(a). One way to achieve proper gate voltages is by sweeping certain gates, which we call screening gates, [S1, S2, S3, and S4 in figure 1(a)] until the currents begin flowing in the correct channels. The bright bars at the edges of the images in figure 1(b) and (d), called walls, are due to the current flowing strictly under those gates. The region with the ridged pattern figure 1(b) is the area of interest – it indicates that the current is not flowing under any screening gates, but instead, it is flowing between them. If there is no current in this region, as shown in figure 1(c), it is necessary to increase the voltage of gates over the current channel to accumulate more electrons in the 2D electron gas and try again.
@@
In gate-defined QD devices, the area of interest usually takes the form of a triangle – thus the name triangle plots – and indicates the formation of an isolated current channel. The patterned texture inside of this region is due to charge defects and variability in gate uniformity near the current channel [28]. While researchers can easily visually identify these regions, finding them automatically using simple threshold analysis is challenging, largely due to experimental variations for different voltage configurations and when measuring different pairs of gates. In moving towards automating the tuning process, it is desirable to have an algorithm that can (1) predict whether the triangle plot indicates a well-behaving current channel and (2) provide an explanation for this prediction.
@@ 2.2 Triangle plots dataset
The triangle plots form a dataset with image data that has a great variety of scales and orientations. We use 902 experimentally acquired triangle plots, with sizes ranging from $31\times 31$ pixels to $171\times 151$ pixels. All images are resized to $64\times 64$ pixels, utilizing a bicubic interpolation provided by the Pillow image library  [29]. This is done so that we can use fixed-parameter filterbanks and synthetic triangles have comparable parameter ranges. Each image is then vectorized following the procedure described in the Gaber filterbank approach (section 2.3) and the synthetic triangle fitting approach (section 2.5).
For the purpose of our study, images that contain the triangle region, as shown in figure 1(a), are considered good. Images that do not contain the triangle region are considered bad. There are a total of 210 good and 692 bad images. The training is performed using five six-fold stratified cross-validations, with $10~{}\%$ of the data withheld for testing purposes.
@@ 2.3 Gabor filterbank approach
To vectorize the image data, we utilize the 2D Gabor wavelet transform, an oriented multi-scale representation that is shown in Ref. [30] to be a model for complex neurons in mammalian vision. Gabor wavelet transform has seen frequent use in numerous computer vision tasks, and serves as an economical and effective feature transform. As such, this representation seems ideal for extracting oriented features, as well as textures, from image data.
Figure 2: (a) A sample experimentally acquired scan, the same as in figure 1(b). (b) The Gabor filters for (a) weighted by their importance as indicated by the EBM model. (c) Synthetic fit to scan shown in (a). The $\mathcal{H}_{B}$, $\mathcal{V}_{B}$, and $\mathcal{D}_{B}$ are the horizontal, vertical, and diagonal optimal boundaries, respectively, and the $\theta$ indicates the orientation of the diagonal boundary.
Definition 1
@@
In practice, we attempt to construct the set of parameters $P$ such that the Fourier transform of the filterbank $G_{P}$ supports prominent frequencies observed in the Fourier transform of the image $u$. This can be accomplished with optimization or other filterbank construction techniques. [31, 32, 33]
To produce the vector representation of the image, we take advantage of heuristics. In particular, the main factor allowing to discriminate good and bad triangles is the presence of the pattern of line-like features at approximately $45^{\circ}$ orientation corresponding to the interaction between the two gates, see figure 1(b). The image is vectorized by taking the $L^{2}$-norm of each filter in the Gabor filterbank with this orientation and at different scales. This achieves a single numerical measure of the filter’s response to the given image, which is ultimately descriptive due to the images’ simple structural content. The specific scales used were obtained through an iterative process. At each iteration, an EBM model is trained on an available filterbank, starting with one that consists of a large number of possible scales and is refined by eliminating scales that are not discriminative. The process is repeated until no further reduction of the filterbank size is observed. The resulting filterbank contains scales determined to be highly discriminative. In our case, the final filterbank consists of six filters at scales $\sigma_{x}=\sigma_{y}=4,8,16$, and orientation $45^{\circ}$ and $-45^{\circ}$ degrees. We also perform additional feature engineering in the form of extracting the estimated location of narrow, edge-like Gabor filters to localize the boundaries of the walls and triangle regions. This technique is used to absorb important spatial features for later use in classification.
Figure 2(b) shows a weighted sum of the filters of the final filterbank with weights defined by the contribution of each term in the EBM. While the resulting filterbank is clearly an incomplete representation of the data, it covers primary discriminant support in the frequency domain between classes. The features are designed to measure the extent of response of the diagonal component of the triangle plot at different scales, where we operate under the intuition that good triangles will have a greater associated response at the ±$\pm 45^{\circ}$ orientations than bad triangles, within a region resembling a fringed isosceles right triangle. The location of the edge-like filters acts as a further sieve for the presence and quality of a triangle region.
@@ 2.4 Synthetic triangle plots
$s_{m,r,b}(x)=\frac{m}{1+{\rm e}^{r(x-b)}},$ (3)
where $m$, $r$, and $b$ control the magnitude, rate, and shift of the sigmoid, respectively. The motivation for this choice is the resemblance of the idealized walls and triangle region to the step function. In practice, the transition from no signal to wall and triangle should be smooth. Triangle plots are assumed to have a horizontal and/or vertical wall corresponding to individual gates, as well as a possible third diagonal region linking the walls. To construct the 2D triangle plots, we compose three sigmoids:
$\Delta_{(x,y)}=\max\big{(}s_{h}(x),s_{v}(y),s_{d}(x\sin\theta+y\cos\theta)\big{)}$ (4)
where $s_{h}$, $s_{v}$, and $s_{d}$ are the parameters for the horizontal, vertical, and diagonal sigmoid, respectively, with $\theta$ indicating the orientation of the diagonal sigmoid. In practice, we evaluate this function on an evenly spaced grid of $(x,y)$ points.
The criteria used to determine the presence of a triangle region are derived from the definition of the triangle plot: we require that $s_{d}(x\sin(\theta)+y\cos(\theta))$ is greater than $s_{h}$ and $s_{v}$. We also impose a constraint that the orientation of the diagonal sigmoid is sufficiently diagonal, with the cutoff defined as $0\leq\theta\leq\nicefrac{{\pi}}{{2}}$.
It is important to restrict the parameters used to define the synthetic triangles to physically reasonable ranges. The simple constraints we chose include a requirement that the sigmoids cannot be infinitely dilated and that the walls must be in the image and have positive heights.
2.5 Synthetic triangle fitting approach

### v5

@@ Abstract
The rapid development of quantum dot (QD) devices for quantum computing has necessitated more efficient and automated methods for device characterization and tuning. Many of the measurements acquired during the tuning process come in the form of images that need to be properly analyzed to guide the subsequent tuning steps. By design, features present in such images capture certain behaviors or states of the measured QD devices. When considered carefully, such features can aid the control and calibration of QD devices. An important example of such images are so-called triangle plots, which visually represent current flow and reveal characteristics important for QD device calibration. While image-based classification tools, such as convolutional neural networks (CNNs), can be used to verify whether a given measurement is good and thus warrants the initiation of the next phase of tuning, they do not provide any insights into how the device should be adjusted in the case of bad images. This is because CNNs sacrifice prediction and model intelligibility for high accuracy. To ameliorate this trade-off, a recent study introduced an image vectorization approach that relies on the Gabor wavelet transform (Schug et al. 2024 Proc. XAI4Sci: Explainable Machine Learning for Sciences Workshop (AAAI 2024) (Vancouver, Canada) pp 1–6). Here we propose an alternative vectorization method that involves mathematical modeling of synthetic triangles to mimic the experimental data. Using explainable boosting machines, we show that this new method offers superior explainability of model prediction without sacrificing accuracy. This work demonstrates the feasibility and advantages of applying explainable machine learning techniques to the analysis of quantum dot measurements, paving the way for further advances in automated and transparent QD device tuning.
@@ 1 Introduction
In many machine learning (ML) applications, there has been a longstanding trade-off between the accuracy and interoperability of candidate models [1, 2, 3]. This is evident in the extreme example of deep neural networks (DNNs), which can offer excellent accuracy for many problems, often surpassing existing methods [4, 5, 6]. Yet, the best-performing ML models are limited in their interpretability due to the number of inaccessible layers. Alternatively, simple techniques like linear models or decision trees allow the user to fully comprehend the internal weights. However, these techniques often cannot model the complex relationships seen in modern datasets. For image data, there has been considerable progress toward finding a middle ground, typically through explaining complex models with surrogates such as the Local Interpretable Model-agnostic Explanations (LIME) [7] and Shapley [8].
@@
In contrast with many black-box ML models, explainable boosting machines (EBMs) are a glass-box method that enables the model to be directly interpretable rather than relying on surrogate explanations [9]. Specifically, EBMs extend generalized additive models to include pairwise interactions, allowing one to observe the relationship between features. EBMs often provide accuracy on par with many black-box models with the additional advantage of enhanced intelligibility, which makes them an appealing replacement for other models, especially in applications of consequence, such as medicine [10, 11, 12] or finance [13]. However, to date, EBMs have not been adapted to any other data type than tabular data.
@@
For spatial data, such as images, interpretability is more challenging. This is partly due to their composition by structures highly correlated at multiple scales in the two-dimensional (2D) space. This feature is one of the main reasons why convolutional neural networks (CNNs) [4] and, more recently, vision transformers (ViTs) [14] have quickly become the dominant ML approach for many computer vision tasks. However, the black-box nature of CNNs and ViTs makes their use prohibitive in applications where a solid understanding of model predictions is necessary. While making CNNs and ViTs more interpretable has been a very active area of research, techniques proposed to date vary in utility in many applications, especially as the depth of the neural network increases [15, 16, 17].
@@
Our previous work developed a methodology that addresses some of these interpretability concerns for measurement data by combining image vectorization with EBMs. Using EBMs as models for image data poses numerous challenges, the principal of which is the mapping from images to a vector representation that could be used directly with EBMs. To achieve this goal, we used the Gabor Wavelet transform and a constrained optimization procedure to extract key image features from the data. We also applied custom feature engineering to tailor this process to the particular dataset [18]. To ensure that the resulting model produces human-agreeable interpretations, we relied on domain knowledge and understanding of the physical systems under investigation to inform the feature extraction process. Here, we demonstrate that the same approach can be successfully applied to assist in the tune-up of accumulation mode Si/SixGe1-x quantum dot (QD) devices. We also propose an alternative image vectorization method involving the generation of synthetic data to approximate the experimentally acquired scans [19]. We then show that both methods result in comparable performance, but the latter produces more intuitive and easier to interpret features.
@@
The paper is organized as follows: in section 2, we provide a brief overview of the scientific context of the problem. Section 2.1 describes the problem of tuning QD devices and introduces the concept of triangle plots. Data used for benchmarking is discussed in section 2.2. The two vectorization methods, the Gabor filterbank approach and the synthetic data modeling approach, are described in section 2.3 and section 2.5, respectively. Finally, the EBMs performance on experimental data vectorized using both methods, as well as using a hybrid approach where the dominant Gabor filter is combined with synthetic data, is presented in section 3. We conclude with a discussion of the future direction in section 4.
@@ 2 Background and methods
In this section, we introduce the scientific context of our problem and the data utilized in our experiments. We also present an overview of the signal processing methods used in the Gabor filterbank approach and the synthetic data modeling approach.
@@ 2.1 Quantum dot tuning problem and triangle plots
Arrays of QDs – interconnected islands of electrons confined in a semiconductor heterostructure with unique properties that allow them to act as artificial atoms – are a leading candidate for use as qubits, the fundamental information carriers in quantum computers [20]. We use a type of gate-defined QD device, as shown in figure 1(a), in which three layers of overlapping gates fabricated on top of a Si/SixGe1-x heterostructure are used to form and control QDs [21, 22]. Scaling these systems to large arrays suitable for quantum computations is a challenging task, as with the number of QDs, the number of gates needed to control them grows, making the manual tuning process unfeasible. An autotuning framework incorporating ML tools was originally proposed and validated off-line in Ref. [23] using premeasured experimental scans capturing a large range of gate voltages and then deployed online (i.e., in situ) to tune a double QD in real-time in Ref. [24]. A detailed description of the tuning process is available in Ref. [25].
@@
Figure 1: (a) False-colored SEM micrograph of a typical Si/SixGe1-x heterostructure device. The red highlighted gates are the screening gates that are swept for the triangle plot. This quad-QD device has three distinct current channels between the screening gates marked with white arrows. One of the channels (in the lower half of the device) contains four QDs in series while the other two channels (in the upper half of the device) are single QD channels for charge-based single-electron-transistor readout. (b) An example of a good triangle plot with high visibility in the triangle region above the background and (c) a corresponding synthetic triangle plot. (c) An example of a bad triangle plot with little to no current in the triangle region and (d) a corresponding synthetic triangle plot.
@@
One sub-problem within the tuning procedure is determining the voltage placement of gates to allow the formation of isolated current channels inside the 2D electron gas formed at the intersection of the Si and SixGe1-x layers in the heterostructure, see figure 1(a). One way to achieve proper gate voltages is by sweeping certain gates, which we call screening gates, [S1, S2, S3, and S4 in figure 1(a)] until the currents begin flowing in the correct channels. The bright bars at the edges of the images in figure 1(b) and (c), called walls, are due to the current flowing strictly under those gates. The region with the ridged pattern figure 1(b) is the area of interest – it indicates that the current is not flowing under any screening gates, but instead, it is flowing between them. If there is no current in this region, as shown in figure 1(c), it is necessary to increase the voltage of gates over the current channel to accumulate more electrons in the 2D electron gas and try again.
@@
In gate-defined QD devices, the area of interest usually takes the form of a triangle – thus the name triangle plots – and indicates the formation of an isolated current channel. Due to the physics principles at play, the only possible orientation of the triangle plot is the one with the right angle in the top-right corner of the image. The patterned texture inside this region is due to charge defects and variability in gate uniformity near the current channel [26]. Combining this with the fact that there is no complete analytical form for the triangle plot at this time makes physical modeling of the measurement impossible. While researchers can easily visually identify these regions, finding them automatically using simple threshold analysis is challenging; largely due to experimental variations for different voltage configurations and when measuring different pairs of gates. In moving towards automating the tuning process, it is desirable to have an algorithm that can (1) predict whether the triangle plot indicates a well-behaving current channel and (2) explain this prediction.
@@ 2.2 Triangle plots dataset
The images in the triangle plots dataset represent various scales and orientations of image data as well as variations in the triangle region size. About $90~{}\%$ of the images were taken at the high temperature of $1.3$ $\mathrm{K}$, instead of roughly $100$ $\mathrm{mK}$, which means that the Coulomb blockade and other gate turn-ons are much less steep than in typical measurements. Since the high-intensity features (e.g., the Coulomb blockade fringes), while prominent to the eye, are less predictive of correct device performance, we focus on the larger-scale features, such as the presence or absence of the triangle region.
For our study, images that contain the triangle region, as shown in figure 1(a), are considered good. Images that do not contain the triangle region are considered bad. We use 902 experimentally acquired triangle plots, of which a total of 210 are labeled as having good triangles and 692 are labeled as having bad triangles.
The images range in size from $31\times 31$ pixels to $171\times 151$ pixels. Prior to analysis, all images are resized to $64\times 64$ pixels, utilizing a bicubic interpolation provided by the Pillow image library [27]. This is done to enable using fixed-parameter filterbanks and synthetic triangles with comparable parameter ranges. While bicubic rescaling can distort the angle of the diagonal region, the majority of the samples that needed preprocessing were in the bad class and thus had minimal to no diagonal activity. In the good class, the maximum observed aspect distortion is within about $13~{}\%$, which corresponds to about $3.5^{\circ}$ difference and does not present significant errors in the context of the Gabor filters. Each image is then vectorized following the procedure described in the Gaber filterbank approach (section 2.3) and the synthetic data modeling approach (section 2.5). The QFlow Triangles dataset containing all experimental images as well as the corresponding synthetic data and Gaber filterbanks is available at Zenodo [28].
@@ 2.3 Gabor filterbank approach
To vectorize the image data, we utilize the 2D Gabor wavelet transform, an oriented multi-scale representation that is shown in Ref. [29] to be a model for complex neurons in mammalian vision. Gabor wavelet transform has seen frequent use in numerous computer vision tasks and serves as an economical and effective feature transform. As such, this representation seems ideal for extracting oriented features, as well as textures, from image data.
Definition.
@@
In practice, we attempt to construct the set of parameters $P$ such that the Fourier transform of the filterbank $G_{P}$ supports prominent frequencies observed in the Fourier transform of the image $u$. This can be accomplished with optimization or other filterbank construction techniques [30, 31, 32].
To produce the vector representation of the image, we take advantage of heuristics. In particular, the main factor allowing the discrimination of good and bad triangles is the presence of the pattern of line-like features at approximately $45^{\circ}$ orientation corresponding to the interaction between the two gates, see figure 1(b). The image is vectorized by taking the $L^{2}$-norm of each filter in the Gabor filterbank with this orientation and at different scales. This achieves a single numerical measure of the filter’s response to the given image, which is ultimately descriptive due to the images’ simple structural content. The specific scales used were obtained through an iterative process. At each iteration, an EBM model is trained on an available filterbank, starting with one that consists of a large number of possible scales and is refined by eliminating scales that are not discriminative. The process is repeated until no further reduction of the filterbank size is observed. The resulting filterbank contains scales determined to be highly discriminative.
In our case, the final filterbank consists of six filters at scales $\sigma_{x}=\sigma_{y}=4,8,16$, and orientation $45^{\circ}$ and $-45^{\circ}$ degrees, where $\lambda$ is fixed to $1$. We also perform additional feature engineering in the form of extracting the estimated location of narrow, edge-like Gabor filters to localize the boundaries of the walls and triangle regions. This technique is used to absorb important spatial features for later use in classification. While other techniques for vectorizing the data based on orientation or scale are possible and, in some instances, advantageous, we chose Gabor because it enables encapsulating the specific diagonal behavior in very few filters. Such vectorization is desirable to ensure that the prediction interpretation is given in terms of relevant features.
Figure 2(b) shows a weighted sum of the filters of the final filterbank for a scan shown in figure 2(a), with weights defined by the contribution of each term in the EBM. While the resulting filterbank is an incomplete representation of the data, it covers primary discriminant support in the frequency domain between classes. The features are designed to measure the extent of response of the diagonal component of the triangle plot at different scales, where we operate under the intuition that good triangles will have a greater associated response at the ±$\pm 45^{\circ}$ orientations than bad triangles, within a region resembling a fringed isosceles right triangle. The location of the edge-like filters acts as a further sieve for the presence and quality of a triangle region. While other vectorization techniques might provide more complete representations of the data and, in some cases, offer superior classification performance, this effect can be diminishing while harming the overall interpretability.
@@ 2.4 Synthetic triangle plots
Figure 2: (a) A sample experimentally acquired scan, the same as in figure 1(b). (b) The Gabor filters for the scan shown in (a), weighted by their importance as indicated by the EBM model. (c) Synthetic fit to scan shown in (a). The $\mathcal{H}_{B}$, $\mathcal{V}_{B}$, and $\mathcal{D}_{B}$ are the horizontal, vertical, and diagonal optimal boundaries, respectively, and the $\theta$ indicates the orientation of the diagonal boundary.
@@
$s_{\rm m,r,b}(x)=\frac{\rm m}{1+{\rm e}^{{\rm r}(x-{\rm b})}},$ (3)
where ${\rm m}$, ${\rm r}$, and ${\rm b}$ control the magnitude, rate, and shift of the sigmoid, respectively. The motivation for this choice is the resemblance of the idealized walls and triangle region to the step function. In practice, the transition from no signal to wall and triangle should be smooth. Triangle plots are assumed to have a horizontal and/or vertical wall corresponding to individual gates, as well as a possible third diagonal region linking the walls. To construct the 2D triangle plots, we compose three sigmoids:
$\Delta_{(x,y)}=\max\big{(}s_{\rm h}(x),s_{\rm v}(y),s_{\rm d}(x\sin\theta+y\cos\theta)\big{)}$ (4)
where $s_{\rm h}$, $s_{\rm v}$, and $s_{\rm d}$ are the parameters for the horizontal, vertical, and diagonal sigmoid, respectively, with $\theta$ indicating the orientation of the diagonal sigmoid. In practice, we evaluate this function on an evenly spaced grid of $(x,y)$ points.
It is important to restrict the parameters used to define the synthetic triangles to physically reasonable ranges. The simple constraints we chose include a requirement that the sigmoids cannot be infinitely dilated and that the walls must be in the image and have positive heights. The criteria used to determine the presence of a triangle region are derived from the definition of the triangle plot: we require that $s_{\rm d}(x\sin(\theta)+y\cos(\theta))$ is greater than $s_{\rm h}$ and $s_{\rm v}$. We also impose a constraint that the orientation of the diagonal sigmoid is sufficiently diagonal, with the cutoff defined as $0\leq\theta\leq\nicefrac{{\pi}}{{2}}$.
2.5 Synthetic data modeling approach